OpenStudy (anonymous):

How do I solve this? -8x + 6 + 4x ≥ 14 .... steps included please.

OpenStudy (anonymous):

Combine like x terms on the left. Move all constants to the right.

OpenStudy (anonymous):

-8x + 4x >= 14-6 = 8 so -4x >= 8 so x >= -8/4 = -2 so x >= -2

OpenStudy (anonymous):

almost, not quite. when you multiply by a negative number you have to switch the directoin of the inequality.

OpenStudy (anonymous):

lol...bad error! Should be -4x >= 8 so x <= -8/4 = -2 so x <= -2

OpenStudy (anonymous):

Thank y'all both! I really appreciate it. :)

OpenStudy (anonymous):

The negative number we're multiplying by here is -1/4, or we are dividing by -4 either way the direction of the inequality should be reversed.